OpenStudy (logan13):

What property is shown in the equation? 7ab • 1 = 7ab A. Associative Property of Multiplication B. Commutative Property of Multiplication C. Identity Property of Multiplication D. Reciprocal Property of Multiplication

OpenStudy (akashdeepdeb):

1 + 0 = 1 This is known as the Identity property of Addition. 0 is known as the additive identity. 1 * 1 = 1 1 is known as the multiplicative identity. Hence, the property shown is the Identity property of Multiplication.

OpenStudy (anonymous):

or it could be commutative 2*3=3*2

OpenStudy (logan13):

um c

OpenStudy (akashdeepdeb):

If it was commutative, they'd surely have given 7ab * 1 in the RHS.

OpenStudy (anonymous):

what is the dot?

OpenStudy (logan13):

multiply

OpenStudy (akashdeepdeb):

Multiplication.